OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (3, 6) and (1, 2)? Check all that apply, there's more than one answer . A.y - 2 = 2(x + 1) B.y - 2 = 2(x - 1) C.y - 6 = 2(x - 3) D.y - 6 = 1/2(x + 3) E.y - 6 = 1/2(x - 3) F.y - 2 = 1/2(x - 1)

Directrix (directrix):

Use the slope formula to get m in y = mx + b. m = (6-2) /( 3-1) = 4/2 = 2/1 or just 2. Think of slope as rise over run.